$500 is deposited in an account which bears 12% p.a. interest compounded monthly. How many years it would take to accumulate $500 to $806.11

1
Expert's answer
2022-03-28T07:45:40-0400

A = P(1+ r/n)nt

806.11 = 500(1 +12/1200)12t

806.11/500 = (1.01)12t

log1.61222 = 12t log 1.01

12t=log 1.6122/log 1.01

12t = 47.9996

t = 47.9996/12

t = 3.9999

4 years